OpenStudy (anonymous):

How can I fin thelimit for + infinity of n*sin(1/n)*cos (1/n^2) ?

hartnn (hartnn):

put n = 1/x in the function. 1/n = ... ? 1/n^2 = ... ?

OpenStudy (anonymous):

if I do that I'll get (1/x) *sin(x)* cos(x^2) right? that will still let me with out knowing it, because (1/x) has the limit of 0, and the other ones have no limit altough they are limited.

hartnn (hartnn):

do you know that the standard limit, \(\Large \lim \limits_{x\rightarrow0} \dfrac{\sin x}{x}=1\)

hartnn (hartnn):

and you can just plug in x=0 in cos (x^2)

OpenStudy (anonymous):

Ah yah, I forgot that I would also have to change the limit from infinit to 0. and i forgot the limit. Thank you very much

hartnn (hartnn):

welcome ^_^